OpenStudy (anonymous):

There are n teams in a league and the number of games played is given by the polynomial n^2 - n=N. A woman's softball league has 6 teams. What is the total number of games to be played?

OpenStudy (anonymous):

plug 6 into the equation. I'm assuming N is the total number of games played?

OpenStudy (chaise):

n^2-n = N 6^2-6=N 36-6=N 30 = N The number of games played is 30.

OpenStudy (anonymous):

Thank you both!